OpenStudy (anonymous):

When combining functions such as f(g(h(x))), which operation should be performed first

OpenStudy (harsimran_hs4):

first you will substitute h(x) suppose f(x) = x^2 g(x) = 3x and h(x) = x+2 so f(g(h(x))) = f(g(x+2)) = f(3(x+2)) = (3(x+2))^2 clear?